Two bodies of masses m and 4m are moving with equal kinetic energies . the ratio of their linear momenta is:
(a) 1:4
(b)4:1
(c)1:2
(d)1:2½

Grade:Upto college level

1 Answers

Deeksha
6 years ago
Kinetic energy = P^2/2mGiven, KE = KE`m = mm` = 4mAccording to question,P^2/2m=P`^2/2(4m)P^2/P`^2=1/4P/P`= √1/4P/P`=1/2